# Classport: Designing Runtime Dependency Introspection for Java

### v1

@@ 5. Experimental Results
5.1. Results for RQ1
Table 2. RQ1 – Static Completeness and Overhead of Embedding Process
@@ 5.1. Results for RQ1
Dependencies Classes Time Space
PDFBox-app 12/12 7,914/7,914 +18.86s 12%
Certificate-ripper 4/5 426/426 +2.68s 14%
mcs 4/4 557/557 +4.43s 10%
batik 6/6 3,679/3,679 +1.15s 29%
checkstyle 34/34 10,691/10,691 +6.18s 17%
zxing 4/4 717/717 +0.51s 11%
@@
The results of RQ1 are summarized in Table 2. It presents the effectiveness of Classport in embedding dependency metadata into Java class files across the open-source projects tested. Each row of the table represents the results of a project on our dataset. The first column reports the completeness of the embedding process considering two aspects. The first one is the number of dependencies embedded compared to the total number of dependencies in the project, and the second one is the number of class files containing the Classport annotation, compared to the total number of class files in the uber-JAR. Then, the table reports the overhead introduced during the build phase by the Embedder. In detail, it reports the absolute time overhead, i.e., the time added by the embedding process to the build, and the space overhead, i.e., the percentage increase in the size of the uber-JAR after embedding.
@@
Consider the first row of the table, which corresponds to the PDFBox project. Classport successfully embeds all 12 dependencies of PDFBox-app, which is the total number of dependencies required by the application. The output uber-JAR contains 7,914 class files, and all of them are annotated with the dependency metadata. The embedding process introduces an absolute time overhead of 18.86 seconds and a space overhead of 12% in the final uber-JAR.
Classport successfully embeds all dependencies and class files in the projects we tested. Thus, 5 out of 6 projects are completely embedded, while Certificate-ripper has one dependency that is not embedded. This is because the dependency is only required for compilation and not for runtime, which is a characteristic known at build time by Maven, and hence it is not included in the uber-JAR. Moreover, classes of uber-JARs of all 6 projects, including the partially embedded project, are annotated with the dependency metadata. This confirms that the embedding process is complete in terms of adding dependency metadata.
@@
The embedding process introduces a moderate performance cost during the build phase. The build time overhead ranges from 0.51 to 18.86 seconds of added time. The value is computed as the time required for the embedding plugin to be run. Considering the entire build process, this time can be negligible in the context of longer testing or packaging steps (Cornejo et al., 2024). The space overhead also varies from 10% to 29% which is considered acceptable given the introduction of a novel feature - visibility into runtime dependencies.
The space overhead depends upon the number of dependencies and identifier names in the core binary content. Certificate-ripper and mcs have about the same number of classes and the exact same number of dependencies embedded, yet the overhead is larger for Certificate-ripper due to longer identifier names. Moreover, the dimension of the class files has a role in the space overhead: since it is computed as a percentage, applications with small class files present a higher percentage overhead than those with big class files, which is the case of batik.
@@
Answer to RQ1: Our experiments show that Classport completely embeds dependency metadata into Java class files. Our benchmark of six diverse Maven projects reflects real-world usage and complexity. Build time and space overhead are an acceptable trade-off for getting a unique feature absent from the Java stack: visibility into runtime dependencies.
5.2. Results for RQ2
Table 3. RQ2 – Dynamic Correctness and Overhead of Introspector
Project Dynamic correctness Runtime Impact
TP FP TN FN Overhead
Functional
Correctness
PDFBox-app 7 0 5 0 0.74%
Certificate-ripper 4 0 0 0 4.27%
mcs 3 0 1 0 1.22%
batik 2 0 4 0 1.24 %
checkstyle 5 0 29 0 0.81%
zxing 3 0 0 0 4.02%
We now assess to what extent we are able to obtain the dependency information during execution. Table 3 summarizes the effectiveness of Classport in detecting dependencies at runtime across the evaluated applications. The table is divided into two parts: the first reports the results regarding the dynamic correctness, while the second shows the runtime impact. Regarding the correctness, the introspection process is defined as correct if the set of runtime dependencies corresponds to the actually used during the execution of the workload, Section 4.4. Specifically, the first column reports the number of true positives (TP), which is the number of dependencies detected by Classport, that are also present in the ground truth, i.e., the dependencies resulting from the mvn dependency:list command. The second column reports the number of false positives (FP), which represents the dependencies that Classport finds but are not present in the ground truth. The third one is about true negatives (TN), i.e., the dependencies that are present in the ground truth but not detected by our tool because the workload does not trigger them. Finally, the fourth column reports the number of false negatives (FN), i.e., the number of dependencies missed by Classport. The part of the table related to the runtime impact is divided into two columns: one reports the percentage of overhead, and the other indicates whether the functional correctness of the target project is preserved after running Classport. For instance, in the case of mcs, Classport successfully detects three dependencies, which are true positives. The one that it does not detect is a true negative because the workload does not trigger it. Classport introduces an overhead of 1.22% and preserves the functional correctness.
Detected Runtime Dependencies The correctness of the identified runtime dependencies is confirmed for all the tested applications. For certificate-ripper and zxing, the workload covers the entire set of dependencies. For the other applications, not all static dependencies are detected at runtime. To investigate whether this is a false negative, we remove from the runtime classpath the dependencies that are not detected by the workload, as explained in subsection 4.4. For PDFBox and checkstyle, the workloads succeed, confirming that Classport correctly identifies the dependencies covered by the workload.
For mcs and batik, the workload fails because one class of removed dependencies is loaded. Manual investigation of the code reveals that these classes are loaded but not executed. This is due to the internal logic of the JVM that requires loading classes even if they are not used. [^10] https://docs.oracle.com/javase/specs/jvms/se17/html/jvms-5.htm This behavior is due to different reasons; in our case, this is related to the following two cases. The JVM loads classes when they are referenced in the code, such as when an object is assigned to a specific type, requiring the JVM to verify that type for safety, even if no actual methods are invoked. Similarly, classes for declared exceptions are loaded, regardless of whether those exceptions ever occur during execution. It would be possible to modify Classport to also intercept the class loading event, but we aim at only detecting executed code. Overall, Classport correctly detects the dependencies used at runtime.
@@ 5.2. Results for RQ2
Overhead. The overhead during the execution of the workload is negligible or low for all the study subjects. It varies from 0.74% to 4.27% which shows that reporting the runtime dependencies does not affect the performance of the application. The overhead is due to the fact that Classport needs to instrument methods invocation and retrieve the dependency metadata from the class files.
@@
Functional correctness We check that Classport does not break any functionality by running the test suite. For all the applications in the dataset, no functional divergence is observed. This means that the bytecode instrumentation done by Classport has no negative effects on the target application.
@@
Answer to RQ2: Classport correctly introspects dependencies at runtime. It does not affect application behavior. The execution overhead is low for all the study subjects. To our knowledge, Classport is the first approach to retrieve dependencies at runtime in the Java ecosystem.
6. Use Cases of Classport
@@
6.1. Runtime Permissions per Dependency
Limiting what resources (e.g., network, filesystem) software components can access at runtime is relevant for security. In the context of Software Supply Chain security, a promising strategy is to assign permissions per dependency, ensuring that each third-party library is granted only the privileges it strictly requires. This permission enforcement per dependency requires precise identification of dependencies at runtime (Amusuo et al., 2024; Cesarano et al., 2025).
@@ 6.1. Runtime Permissions per Dependency
6.2. Vulnerability Detection at Runtime
Vulnerability detection tools aim to identify known vulnerabilities in third-party libraries. Identifying dependencies that are actually used at runtime is fundamental to avoid false positives or false negatives in vulnerability assessment (Williams et al., 2025).
@@ 7. Related Work
7.1. Embedding Information into Executables and Binaries
Boucher et al. (Boucher and Anderson, 2023) propose a novel approach to identifying software dependencies, introducing ABOM, Automatic Bill Of Material. The main idea of ABOM is to query an executable for its dependencies to facilitate vulnerability detection. ABOM automatically embeds hashes of dependencies into compiled binaries during the build process. These hashes are stored efficiently in a probabilistic data structure called Compressed Bloom Filters. It allows for rapid querying to detect the presence of a specific dependency and hence to identify the vulnerabilities associated with it. However, since the data structure is probabilistic in nature, it can report the presence of a dependency even if it is not actually present. Our approach cannot generate false positives because it directly relies on Maven and a Maven build success implies that all dependencies for the project are resolved.
Seshadi et al. (Seshadri et al., 2024) present OmniBOR, which embeds a content-based identifier called gitoid into executables, including Java class files, to link them to an external Artifact Dependency Graph built during compilation. This approach requires build-time instrumentation and external storage. In contrast, our method embeds Maven GAV metadata directly into class files, allowing immediate runtime access via a Java agent without relying on external resolution. Our technique is dedicated to Java and uniquely supports lightweight, runtime introspection of runtime dependencies.
Quatch et al. (Quach et al., 2018) propose a solution that combines static and dynamic analysis to identify and remove the unused code during the program loading process. Their framework generates function-level dependency graphs and embeds this information into an optional section of the ELF binary. This embedded dependency information is then used by a piece-wise loader at runtime to identify and remove unused code from the program’s memory. Our approach is more general since it brings dependency information from the build-time context to the runtime one and it is not limited to use cases like debloating.
7.2. Runtime Identification of Dependencies
@@
Ponta et al. (Ponta et al., 2020) present a novel code-centric and usage-based method for detecting, assessing, and mitigating vulnerabilities in open-source software dependencies. This work builds upon their previous approaches (Plate et al., 2015; Ponta et al., 2018), and is implemented in the open-source tool Eclipse Steady. [^11] https://github.com/eclipse-steady/steady They use a dynamic call graph of dependencies to check the reachability of the vulnerable code. They instrument the target Java application and search if, during the execution, it reaches the vulnerable code, i.e., if it uses vulnerable methods. The difference with our approach is that they focus on the vulnerable dependencies and not on an overall identification of the actually executed dependencies.
Soto-Valero et al. (Soto-Valero et al., 2023) propose a tool to detect bloated code in Java and to remove it by transforming the bytecode of the compiled project. They use four code coverage tools along with probes to identify methods that are not executed during runtime. If no methods in any class are executed within a dependency, the dependency is considered bloated. They only work at testing time. However, we embed dependency information directly in the class file in JARs, making this information available at production time as well.
Amusuo et al. (Amusuo et al., 2024) implement a tool that is able to create a policy file with a mapping of required permissions for every dependency. The policies are then enforced at runtime. To create a policy file, they infer dependency namespaces by assuming that all classes within a JAR share a common root directory path, which is then mapped to the Maven Group and Artifact ID. In contrast, we directly retrieve dependency information at runtime from metadata embedded in the class files, avoiding fragile assumptions based on package name heuristics.
8. Conclusion

### v2

@@ 5 Experimental Results
5.1 Results for RQ1
Table 2: RQ1 – Static Completeness and Overhead of Embedding Process
@@ 5.1 Results for RQ1
Dependencies Classes Time (s) Space (MB)
PDFBox-app (app) 12/12 7,914/7,914 11.52%(2.4) 13.82%(1.9)
Certificate-ripper (itself) 4/5 426/426 6.80%(0.2) 13.17%(0.1)
mcs (itself) 4/4 557/557 7.62%(0.3) 9.88%(0.1)
BioJava (biojava-aa-prop) 31/32 7,629/7,629 7.17%(1.2) 17.19%(2.4)
Checkstyle (itself) 34/37 10,691/10,691 11.38%(2.7) 17.20%(3.4)
GraphHopper (graphhoper-web) 122/166 20,965/20,965 16.80%(6.5) 12.96%(6.1)
@@
The results of RQ1 are summarized in Table 2. It presents the effectiveness of Classport in embedding dependency metadata into Java class files across the open-source projects tested. Each row of the table represents the results of a project on our dataset. The first column lists the projects and their executable uberjar from our dataset described in subsection 4.2. The second column reports the completeness of the embedding process considering two aspects. The first aspect is the number of dependencies embedded compared to the total number of dependencies in the project, and the second one is the number of class files containing the Classport annotation, compared to the total number of class files in the original uber-JAR. Then, the table reports the overhead introduced during the build phase by the Embedder. We report the embedding time overhead, i.e., the time added by the embedding process to the build, and the space overhead, i.e., the size of the uber-JAR after embedding. We also show the absolute increase in paranthesis.
@@
Consider the first row of the table, which corresponds to the PDFBox project. Classport successfully embeds all 12 dependencies of PDFBox-app, which is the total number of dependencies required by the application. The output uber-JAR contains 7,914 class files, and all of them are annotated with the dependency metadata. For this project, the embedding process introduces a time overhead of 11.52% and a space overhead of 13.82% in the final uber-JAR.
There are two projects, PDFBox and mcs, which are completely embedded, while the other projects are partially embedded. There are three reasons why an uber-JAR does not have all the dependencies as declared in the pom.xml file. The first one is the case where a dependency JAR does not contain any class files. It may be empty, which happens for dependency com.google.guava:listenablefuture:9999.0-empty-to-avoid-conflict
-with-guava
, or it may contain files such as JavaScript, Kotlin metadata, or other non-class files. Checkstyle and GraphHopper are examples of projects that have dependencies that do not contain any class files.
The second reason is where a dependency is only necessary for compilation and not for runtime, which is called a provided dependency in Maven. This dependency is not included in the uber-JAR by default. Certificate Ripper and Checkstyle are examples of projects that have dependencies that are only necessary for compilation and not for runtime. Finally, the third reason is where classes of a dependency are shadowed by the other. In this case, BioJava has the dependency javax.xml.bind:jaxb-api which shadows jakarta.xml.bind:jakarta.xml.bind-api. Note that shadowing is not a consequence of the Embedder, but a consequence of how Maven packages the uber-JAR [22]. In all three cases, the classes from those dependencies are not executed. Recall that Classport is designed to detect dependencies that are executed during runtime. Therefore, these dependencies are rightfully not detected by Classport.
@@
The embedding process introduces a moderate performance cost during the build phase. On an absolute scale, the embedding time overhead ranges from 0.2s to 6.5s which we claim is acceptable in a build. The space overhead varies from 0.1MB to 6.1MB which is considered acceptable given the introduction of a novel feature: visibility into runtime dependencies. The space overhead depends upon the number of dependencies and identifier names in the core binary content. Certificate Ripper and mcs have about the same number of classes and the exact same number of dependencies embedded, yet the overhead is larger for Certificate Ripper due to longer identifier names.
@@
Answer to RQ1: Our experiments show that Classport completely embeds dependency metadata into Java class files. Our benchmark of six diverse Maven projects demonstrates that Classport works with real-world and complex software. Build time and space overhead are an acceptable trade-off for getting a unique feature absent from the Java stack: runtime visibility into dependencies.
5.2 Results for RQ2
Table 3: RQ2 – List of Runtime Dependencies and Overhead of Introspector
Project Runtime Dependencies Overhead (s) Functional Correctness
PDFBox 13 4.11%(1.9)
Certificate-ripper 4 2.95%(0.1)
mcs 4 5.96%(0.3)
BioJava 23 8.54%(56.8)
Checkstyle 22 5.28%(8.8)
GraphHopper 104 5.22%(11.0)
We now assess to what extent we are able to obtain the dependency information during execution. Table 3 summarizes the effectiveness of Classport in detecting dependencies at runtime across the evaluated applications. The table is divided into three parts: the first part reports the results regarding the detected runtime dependencies, while the second shows the runtime overhead (as opposed to the build overhead in RQ1), and the third one reports the functional correctness.
Detected Runtime Dependencies This is the number of dependencies detected by Classport during the execution of the workload. For Certificate Ripper and mcs, the workload covers the entire set of embedded dependencies. For the other applications, not all embedded dependencies are detected at runtime. On manual inspection, we found that this is due to two reasons. First, some dependencies are not exercised by the workload or are bloated. Such as the dependency io.dropwizard.logback:logback-throttling-appender:1.5.1 of GraphHopper is not used by the workload. Second, some dependencies have classes that are loaded by the Java Virtual Machine (JVM) but not executed. This happens because JVM loads classes to check for type safety, declared exceptions, and extracting metadata from annotations. For example, the dependency com.google.errorprone:error_prone_annotations:2.36.0 only has annotations. Annotations in Java are not executed as they don’t have any methods. For these two reasons, Classport does not detect these dependencies because it only considers dependencies that are executed during runtime.
@@ 5.2 Results for RQ2
Overhead. The overhead during execution of ilow for all the study subjects. It varies from 2.95% to 8.54%. The overhead is mostly due to the fact that Classport needs to instrument methods invocation with an agent.
@@
Functional correctness Classport does not break any functionality as the entire test suite runs with the same result with or without introspection. This means that the bytecode instrumentation done by Classport does not introduce bugs on the target application.
@@
Answer to RQ2: Classport can successfully introspect dependencies at runtime, and allows applications to identify which dependencies are being currently executed. Classport does not affect application behavior. The execution overhead is low for all study subjects. To our knowledge, Classport is the first system in the Java ecosystem to support dependency introspection at runtime.
6 Use Cases of Classport
@@
6.1 Runtime Permissions per Dependency
Limiting what resources (e.g., network, filesystem) software components can access at runtime is relevant for security. In the context of Software Supply Chain security, a promising strategy is to assign permissions per dependency, ensuring that each third-party library is granted only the privileges it strictly requires. This permission enforcement per dependency requires precise identification of dependencies at runtime [13, 23].
@@ 6.1 Runtime Permissions per Dependency
6.2 Vulnerability Detection at Runtime
Vulnerability detection tools aim to identify known vulnerabilities in third-party libraries. Identifying dependencies that are actually used at runtime is fundamental to avoid false positives or false negatives in vulnerability assessment [24].
@@ 7 Related Work
7.1 Embedding Information into Executables and Binaries
Boucher et al. [25] propose a novel approach to identifying software dependencies, introducing ABOM, Automatic Bill Of Material. The main idea of ABOM is to query an executable for its dependencies to facilitate vulnerability detection. ABOM automatically embeds hashes of dependencies into compiled binaries during the build process. These hashes are stored efficiently in a probabilistic data structure called Compressed Bloom Filters. It allows for rapid querying to detect the presence of a specific dependency and hence to identify the vulnerabilities associated with it. However, since the data structure is probabilistic in nature, it can report the presence of a dependency even if it is not actually present. Our approach cannot generate false positives because it directly relies on Maven and a Maven build success implies that all dependencies for the project are resolved.
Seshadi et al. [26] present OmniBOR, which embeds a content-based identifier called gitoid into executables, including Java class files, to link them to an external Artifact Dependency Graph built during compilation. This approach requires build-time instrumentation and external storage. In contrast, our method embeds Maven GAV metadata directly into class files, allowing immediate runtime access via a Java agent without relying on external resolution. Our technique is dedicated to Java and uniquely supports lightweight, runtime introspection of runtime dependencies.
Quatch et al. [27] propose a solution that combines static and dynamic analysis to identify and remove the unused code during the program loading process. Their framework generates function-level dependency graphs and embeds this information into an optional section of the ELF binary. This embedded dependency information is then used by a piece-wise loader at runtime to identify and remove unused code from the program’s memory. Our approach is more general since it brings dependency information from the build-time context to the runtime one and it is not limited to use cases like debloating.
7.2 Runtime Identification of Dependencies
@@
Ponta et al. [28] present a novel code-centric and usage-based method for detecting, assessing, and mitigating vulnerabilities in open-source software dependencies. This work builds upon their previous approaches [29, 30], and is implemented in the open-source tool Eclipse Steady. [^9] https://github.com/eclipse-steady/steady They use a dynamic call graph of dependencies to check the reachability of the vulnerable code. They instrument the target Java application and search if, during the execution, it reaches the vulnerable code, i.e., if it uses vulnerable methods. The difference with our approach is that they focus on the vulnerable dependencies and not on an overall identification of the actually executed dependencies.
Soto-Valero et al. [31] propose a tool to detect bloated code in Java and to remove it by transforming the bytecode of the compiled project. They use four code coverage tools along with probes to identify methods that are not executed during runtime. If no methods in any class are executed within a dependency, the dependency is considered bloated. They only work at testing time. However, we embed dependency information directly in the class file in JARs, making this information available at production time as well.
Amusuo et al. [13] implement a tool that is able to create a policy file with a mapping of required permissions for every dependency. The policies are then enforced at runtime. To create a policy file, they infer dependency namespaces by assuming that all classes within a JAR share a common root directory path, which is then mapped to the Maven Group and Artifact ID. In contrast, we directly retrieve dependency information at runtime from metadata embedded in the class files, avoiding fragile assumptions based on package name heuristics.
8 Conclusion